# Cohomogeneity one $\mathsf{RCD}$-spaces

### v2

@@ 3.2. Group Diagram
Let $G$ be a compact Lie group, acting by isometries on a geodesic space $(X,d)$, such that $X^{\ast}$ is isometric to $[-1,1]$. Given $t_{0}\in(-1,1)$ there exists $\varepsilon>0$ and a geodesic $\gamma\colon(-\varepsilon+t_{0},\varepsilon+t_{0})\to X$ with $\gamma(t_{0})\in\pi^{-1}(t_{0})$ realizing the distance between the orbits. That is, for any $t,s\in(-\varepsilon+t_{0},\varepsilon+t_{0})$ we have $d(G(\gamma(s)),G(\gamma(t)))=d(\gamma(s),\gamma(t))$.
Proof.
We fix $x_{0}\in\pi^{-1}(t_{0})=G(x_{0})$, and choose $\varepsilon>0$ small enough so that for the slice $S_{x_{0}}$ through $x_{0}$ we have $(-\varepsilon+t_{0},\varepsilon+t_{0})\subset\pi(S_{x_{0}})$. We consider the minimal geodesic $\gamma^{\ast}\colon(-\varepsilon+t_{0},\varepsilon+t_{0})\to[-1,1]$ with $\gamma^{\ast}(t_{0})=t_{0}$, and fix $t_{0}-\varepsilon<t_{0}-s\leq t_{0}<t_{0}+t<t_{0}+\varepsilon$. Then there exist points $y,z\in S_{x_{0}}$ such that $y\in S_{x_{0}}\cap\pi^{-1}(t_{0}-s)$ and $z\in S_{x_{0}}\cap\pi^{-1}(t_{0}+t)$. Moreover, by construction of $S_{x_{0}}$, there exist minimizing geodesics $\tilde{\gamma}_{1}\colon[-s,0]\to X$ and $\tilde{\gamma}_{1}\colon[0,t]\to X$ such that $\tilde{\gamma}_{i}(r)\in S_{x_{0}}$ for $r\in[-s,0]$ if $i=1$ and $r\in[0,t]$ if $i=2$. This implies that for the concatenation of $\tilde{\gamma}_{2}$ and $\tilde{\gamma}_{1}$ we have by construction $\pi(\tilde{\gamma}_{2}\ast\tilde{\gamma}_{1})=\gamma^{\ast}|_{[t_{0}-s,t_{0}+t]}$, and moreover $d(\tilde{\gamma}_{1}(-r+t_{0}),x_{0})=r=d^{\ast}(\gamma^{\ast}(t_{0}-r),t_{0})$ for $r\in[0,s]$, and $d(\tilde{\gamma}_{2}(\ell+t_{0}),x_{0})=\ell=d^{\ast}(\gamma^{\ast}(t_{0}+\ell),t_{0})$ for $\ell\in[0,t]$. From this we have for $t_{0}-s\leq t_{0}-r\leq t_{0}<t_{0}+\ell\leq t_{0}+t$:
$d(\tilde{\gamma}_{1}(t_{0}-r),\tilde{\gamma}_{2}(t_{0}+\ell))$ $\leq d(\tilde{\gamma}_{1}(t_{0}-r),x_{0})+d(x_{0},\tilde{\gamma}_{2}(t_{0}+\ell))$
$d^{\ast}(\gamma^{\ast}(t_{0}-r),t_{0})+d^{\ast}(t_{0},\gamma^{\ast}_{2}(t_{0}+\ell))$
$=d^{\ast}(\gamma^{\ast}(t_{0}-r),\gamma^{\ast}_{2}(t_{0}+\ell))$
$=d(G(\tilde{\gamma}_{1}(t_{0}-r)),G(\tilde{\gamma}_{2}(t_{0}+\ell)))$
$\leq d(\tilde{\gamma}_{1}(t_{0}-r),\tilde{\gamma}_{2}(t_{0}+\ell)).$
Thus we conclude that for the concatenation $\gamma:=\tilde{\gamma}_{2}\ast\tilde{\gamma}_{1}\colon[t_{0}-s,t_{0}+t]$ we have for all $r,\ell\in[t_{0}-s,t_{0}+t]$ that
$\mathrm{dist}(\gamma(r),\gamma(\ell))=|r-\ell|.$
Therefore, it follows that $\gamma$ is a geodesic realizing the distance between the orbits. ∎
Lemma 3.13.
Let $G$ be a compact Lie group acting by isometries on a geodesic space $(X,d)$, such that the orbit space $X^{\ast}$ is isometric to $[-1,1]$. Then there exists a minimal geodesic $\gamma\colon[-1,1]\to X$ realizing the distance between the orbits. That is, for any $t,s\in[-1,1]$ we have $d(G(\gamma(s)),G(\gamma(t)))=d(\gamma(s),\gamma(t))$.
Proof.
We fix $\gamma^{\ast}\colon(-1,1)\to(-1,1)$ the minimal geodesic with $\gamma^{\ast}(0)=0$ consider a cover $\{U_{i}\}_{i\in\Lambda}$ of $(-1,1)$ by sets satisfying the hypothesis of Lemma 3.12. Now fix $x_{0}\in\pi^{-1}(0)$, and assume that $0\in U_{i_{0}}$, and $U_{j_{0}}\cap U_{i_{0}}\neq\varnothing$, $U_{k_{0}}\cap U_{i_{0}}\neq\varnothing$. We consider the lift $\gamma_{i_{0}}$ of $\gamma^{\ast}|_{U_{i_{0}}}$ through $x_{0}$ given by Lemma 3.12. We observe that we can extend this geodesic to $U_{j_{0}}$ and $U_{k_{0}}$, since the end points of $\gamma_{i_{0}}$ determine points in $\pi^{-1}(U_{j_{0}})$ and $\pi^{-1}(U_{k_{0}})$, and we can apply again Lemma 3.12. Continuing in this fashion we obtain a geodesic $\gamma\colon(-1,1)\to(X,d)$ which is a lift of $\gamma^{\ast}$. Moreover, by construction for $s,t\in(-1,1)$ we have
$d^{\ast}(\gamma^{\ast}(s),\gamma^{\ast}(t))=d(G(\gamma(s)),G(\gamma(t)))=d(\gamma(s),\gamma(t))$
To finish we observe that by [76, Theorem 1] (see also [7, Theorem 1A], [8]) we can extend $\gamma$ to a continuous curve $\gamma\colon[-1,1]\to(X,d)$. Moreover, by continuity of the quotient map $\pi$, and the distances $d^{\ast}$ and $d$ we have
$d(G(\gamma(-1)),G(\gamma(1)))$ $=d^{\ast}(\gamma^{\ast}(-1),\gamma^{\ast}(1))$
$=d^{\ast}(\pi(\gamma)(-1),\pi(\gamma)(1))$
$=\lim_{\begin{subarray}{c}s\to-1\\ t\to 1\end{subarray}}d^{\ast}(\pi(\gamma)(s),\pi(\gamma)(t))$
$=\lim_{\begin{subarray}{c}s\to-1\\ t\to 1\end{subarray}}d(G(\gamma(s)),G(\gamma(t)))$
$=\lim_{\begin{subarray}{c}s\to-1\\ t\to 1\end{subarray}}d(\gamma(s),\gamma(t))$
$=d(\gamma(-1),\gamma(1)).$
This proves our claim. ∎
With this we can relate the isotropy groups as follows.
Theorem 3.14.
Let $(X,d)$ be a geodesic space, and $G$ a compact Lie group acting by isometries, such that $(X^{\ast},d^{\ast})$ is isometric to $[-1,1]$. Then we have three isotropy subgroups $H$, $G_{-}$, and $G_{+}$, corresponding to orbits in the interior $(-1,1)$, and the end points $\{-1\}$ and $\{+1\}$ respectively. Moreover, $H$ is a subgroup of both $G_{-}$ and $G_{+}$.
Proof.
We fix $\gamma\colon[-1,1]\to X$ a geodesic realizing the distance between the orbits, guaranteed to exist by Lemma 3.13. Then by Lemma 3.10 the conclusions follow, by taking $H:=G_{\gamma(0)}$, and $G_{\pm}=G_{\gamma(\pm)}$. ∎
The tuple $(G,H,G_{+},G_{-})$ obtained in Theorem 3.14 depends a priori on a choice of a curve. In the following lemma we show how two tuples associated to two different actions relate to each other.
Lemma 3.15.
Let $(X,d)$ be a geodesic space, and $G$ a compact Lie group acting by isometries, such that $(X^{\ast},d^{\ast})$ is isometric to $[-1,1]$. Let $\gamma_{1}$ and $\gamma_{2}$ be two geodesics realizing the distance between the orbits, and $(G,H^{1},K^{1}_{+},K^{1}_{-})$ and $(G,H^{2},K^{2}_{+},K^{2}_{-})$ their associated tuples. Then for some $g_{0}\in G$ and $a_{\pm}\in N(H^{1})_{0}$ we have
$(G,H^{2},K_{+}^{2},K_{-}^{2})=(G,g_{0}H^{1}g_{0}^{-1},g_{0}a_{+}K_{+}^{2}a_{+}^{1}g_{0}^{-1},g_{0}a_{-}K_{-}^{2}a_{-}^{1}g_{0}^{-1}).$
@@
We consider $\{x_{0}^{1}\}=\pi^{-1}(0)\cap\gamma_{1}([-1,1])$ and $\{x_{0}^{1}\}=\pi^{-1}(0)\cap\gamma_{2}([-1,1])$. Then there exists $g_{0}\in G$, such that $gx^{1}_{0}=x_{0}^{2}$. Then $G_{x_{0}^{2}}=g_{0}G_{x^{1}_{0}}g_{0}^{-1}$. Moreover, $\tilde{\gamma}_{2}:=g_{0}^{-1}\gamma_{2}$ is also a geodesic with $\tilde{\gamma}_{2}(0)=x_{0}^{1}$. Moreover, there exists $a\colon[-1,1]\to G$ such that $a(t)\gamma_{1}(t)=\tilde{\gamma}_{2}(t)$. Observe that $a(0)=e$, and thus we have that $G_{\tilde{\gamma}_{2}(t)}=a(t)G_{\gamma_{1}(t)}(a(t))^{-1}$. But by Lemma 3.10 we have that $G_{\gamma_{1}(t)}=\Gamma_{\gamma_{1}(0)}=\Gamma_{x_{0}^{1}}=\Gamma_{\tilde{\gamma}_{2}(0)}=G_{\tilde{\gamma}_{2}(t)}$. Writing $H:=\Gamma_{x_{0}^{1}}$ we conclude that $a(t)\in N(H)$ for all $t\in[-1,1]$. Thus by setting $a_{\pm}:=a(\pm 1)$ we have that
$\tilde{K}_{\pm}^{2}=G_{\tilde{\gamma}_{2}(\pm 1)}=a_{\pm}G_{\gamma_{1}(\pm 1)}a_{\pm}^{-1}=a\pm K_{\pm}^{1}a_{\pm}^{-1}.$
Since $\tilde{\gamma}_{2}(\pm)=g_{0}^{-1}\gamma_{2}(\pm)$, we have
$\tilde{K}_{\pm}^{2}=G_{\tilde{\gamma}_{2}(\pm 1)}=g_{0}^{-1}G_{\gamma_{2}(\pm)}g_{0}=g_{0}^{-1}K_{\pm}^{2}g_{0}.$
@@
Let $G$ be a compact Lie group acting on two geodesic spaces $(X_{1},d_{1})$ and $(X_{2},d_{2})$ by isometries and $X_{1}^{\ast}$ and $X_{2}^{\ast}$ are isometric to $[-1,1]$. Moreover, assume that we have associated the tuples $(G,H^{1},K^{1}_{+},K_{-}^{1})$ and $(G,H^{2},K^{2}_{+},K_{-}^{2})$ respectively. If the spaces are equivariantly homeomorphic then there exists $g_{0}\in G$ and $a_{\pm}\in N(H^{1})_{0}$ such that $H^{2}=g_{0}H^{1}g_{0}^{-1}$ and either $K_{\pm}^{2}=g_{0}a_{\pm}K_{\pm}^{1}a_{\pm}^{-1}g_{0}^{-1}$ or $K_{\pm}^{2}=g_{0}a_{\pm}K_{\mp}^{1}a_{\pm}^{-1}g_{0}^{-1}$.

### v3

@@ 3.1. Cohomogeneity one Group diagrams
Let $(X,d,\mathfrak{m})$ be an essentially non-branching space, and $G$ a compact Lie group acting by isometries, such that $(X^{\ast},d^{\ast})$ is isometric to [1,1][-1,1]. Let $\gamma^{1}$ and $\gamma^{2}$ be two geodesics realizing the distance between the orbits, and $(G,H^{1},K^{1}_{+},K^{1}_{-})$ and $(G,H^{2},K^{2}_{+},K^{2}_{-})$ their associated tuples. Then for some $g_{0}\in G$ and $a_{\pm}\in N(H^{1})_{0}$ we have
$(G,H^{2},K_{+}^{2},K_{-}^{2})=(G,g_{0}H^{1}g_{0}^{-1},g_{0}a_{+}K_{+}^{2}a_{+}^{-1}g_{0}^{-1},g_{0}a_{-}K_{-}^{2}a_{-}^{-1}g_{0}^{-1}).$
@@
We consider $\{x_{0}^{1}\}=\pi^{-1}(1/2)\cap\gamma^{1}([0,1])$ and $\{x_{0}^{2}\}=\pi^{-1}(1/2)\cap\gamma^{2}([0,1])$. Then there exists $g_{0}\in G$, such that $gx^{1}_{0}=x_{0}^{2}$. Then $G_{x_{0}^{2}}=g_{0}G_{x^{1}_{0}}g_{0}^{-1}$. Moreover, $\tilde{\gamma}^{2}:=g_{0}^{-1}\gamma^{2}$ is also a geodesic with $\tilde{\gamma}^{2}(0)=x_{0}^{1}$. Moreover, there exists $a\colon[0,1]\to G$ such that $a_{t}\gamma^{1}_{t}=\tilde{\gamma}^{2}_{t}$. Observe that $a_{1/2}=e$, and thus we have that $G_{\tilde{\gamma}^{2}_{t}}=a_{t}G_{\gamma^{1}_{t}}a_{t}^{-1}$. But by ?THM? LABEL:L:_Kleiners_Lemma we have that $G_{\gamma^{1}_{t}}=G_{\gamma^{1}_{1/2}}=G_{x_{0}^{1}}=G_{\tilde{\gamma}^{2}_{1/2}}=G_{\tilde{\gamma}^{2}_{t}}$ for all $t\in(0,1)$. Writing $H:=G_{x_{0}^{1}}$ we conclude that $a_{t}\in N(H)$ for all $t\in[0,1]$. Thus by setting $a_{\pm}:=a_{\pm 1}$ we have that
$\tilde{K}_{\pm}^{2}=G_{\tilde{\gamma}^{2}_{\pm 1}}=a_{\pm}G_{\gamma^{1}_{\pm 1}}a_{\pm}^{-1}=a_{\pm}K_{\pm}^{1}a_{\pm}^{-1}.$
Since $\tilde{\gamma}^{2}_{\pm}=g_{0}^{-1}\gamma^{2}_{\pm}$, we have
$\tilde{K}_{\pm}^{2}=G_{\tilde{\gamma}^{2}_{\pm 1}}=g_{0}^{-1}G_{\gamma^{2}_{\pm 1}}g_{0}=g_{0}^{-1}K_{\pm}^{2}g_{0}.$
@@
Let $G$ be a compact Lie group acting on two essentially non-branching spaces $(X_{1},d_{1},\mathfrak{m}_{1})$ and $(X_{2},d_{2},\mathfrak{m}_{2})$ by isometries such that $(X_{1}^{\ast},d_{1}^{\ast})$ and $(X_{2}^{\ast},d_{2}^{\ast})$ are isometric to [1,1][-1,1]. Moreover, assume that we have associated the tuples $(G,H^{1},K^{1}_{+},K_{-}^{1})$ and $(G,H^{2},K^{2}_{+},K_{-}^{2})$ respectively. If the spaces are equivariantly homeomorphic then there exists $g_{0}\in G$ and $a_{\pm}\in N(H^{1})_{0}$ such that $H^{2}=g_{0}H^{1}g_{0}^{-1}$ and either $K_{\pm}^{2}=g_{0}a_{\pm}K_{\pm}^{1}a_{\pm}^{-1}g_{0}^{-1}$ or $K_{\pm}^{2}=g_{0}a_{\pm}K_{\mp}^{1}a_{\pm}^{-1}g_{0}^{-1}$.